# Dual-Regularized Riccati Recursions for Interior-Point Optimal Control

### v4

@@ II-B The Regularized Interior Point Method
$\mathcal{A}(x,s,y,z;\mu,\eta)$
$=$ $\mathcal{L}(x,s,y,z;\mu)+\frac{\eta}{2}(\lVert c(x)\rVert^{2}+\lVert g(x)+s\rVert^{2}).$
@@
$\begin{bmatrix}P&0&C^{T}&G^{T}\\ 0&S^{-1}Z&0&I\\ C&0&-\frac{1}{\eta}I&0\\ G&I&0&-\frac{1}{\eta}I\end{bmatrix}\begin{bmatrix}\Delta x\\ \Delta s\\ \Delta y\\ \Delta z\end{bmatrix}$$\nabla\mathcal{L}$, (1)
where $\nabla\mathcal{L}$ is evaluated at $(x,s,y,z;\mu)$, $P$ is a positive-definite approximation of $\nabla^{2}_{xx}\mathcal{L}(x,s,y,z;\mu)$, $C=J(c)(x)$, $G=J(g)(x)$, and $S,Z$ denote the diagonal matrices with entries $s,z$ respectively.
We call $x,s$ the primal variables and $y,z$ the dual variables. As usual, we require that $s,z>0$ always holds. Below, we will show that primal search direction determined by the regularized interior point method is guaranteed to be a descent direction of $\mathcal{A}(\cdot,\cdot,y,z;\mu,\eta)$ at $(x,s)$ unless the primal variables have already converged.
@@
$\begin{bmatrix}P&0&C^{T}&G^{T}\\ 0&S^{-1}Z&0&I\\ C&0&-\frac{1}{\eta}I&0\\ G&I&0&-\frac{1}{\eta}I\end{bmatrix}\begin{bmatrix}\Delta x\\ \Delta s\\ \Delta y-\eta c(x)\\ \Delta z-\eta(g(x)+s)\end{bmatrix}$ (2)
$=$ $-\begin{bmatrix}\nabla_{x}\mathcal{A}(x,s,y,z;\mu,\eta)\\ \nabla_{s}\mathcal{A}(x,s,y,z;\mu,\eta)\\ 0\\ 0\end{bmatrix}$
@@
When $\Delta x\neq 0$ or $\Delta s\neq 0$, $D(\mathcal{A}(\cdot,\cdot,y,z;\mu,\eta);(\Delta x,\Delta s))(x,s)<0$.
@@
$D(\mathcal{A}(\cdot,\cdot,y,z;\mu,\eta);(\Delta x,\Delta s))(x,s)$
$=$ $\begin{bmatrix}\Delta x^{T}&\Delta s^{T}\end{bmatrix}\begin{bmatrix}\nabla_{x}\mathcal{A}(x,s,y,z;\mu,\eta)\\ \nabla_{s}\mathcal{A}(x,s,y,z;\mu,\eta)\\ \end{bmatrix}$
$=$ $\begin{bmatrix}\Delta x^{T}&\Delta s^{T}&0&0\end{bmatrix}\begin{bmatrix}\nabla_{x}\mathcal{A}(x,s,y,z;\mu,\eta)\\ \nabla_{s}\mathcal{A}(x,s,y,z;\mu,\eta)\\ 0\\ 0\end{bmatrix}$
$=$ $-\begin{bmatrix}\Delta x^{T}&\Delta s^{T}&0&0\end{bmatrix}\begin{bmatrix}P&0&C^{T}&G^{T}\\ 0&S^{-1}Z&0&I\\ C&0&-\frac{1}{\eta}I&0\\ G&I&0&-\frac{1}{\eta}I\end{bmatrix}$
$\begin{bmatrix}\Delta x\\ \Delta s\\ \Delta y-\eta c(x)\\ \Delta z-\eta(g(x)+s)\end{bmatrix}$
$=$ $-\begin{bmatrix}\Delta x^{T}&\Delta s^{T}\end{bmatrix}\begin{bmatrix}P&0\\ 0&S^{-1}Z\end{bmatrix}\begin{bmatrix}\Delta x\\ \Delta s\end{bmatrix}$
$-\begin{bmatrix}\Delta x^{T}&\Delta s^{T}\end{bmatrix}\begin{bmatrix}C^{T}&G^{T}\\ 0&I\end{bmatrix}\begin{bmatrix}\Delta y-\eta c(x)\\ \Delta z-\eta(g(x)+s)\end{bmatrix}$
$=$ $-\begin{bmatrix}\Delta x^{T}&\Delta s^{T}\end{bmatrix}\begin{bmatrix}P&0\\ 0&S^{-1}Z\end{bmatrix}\begin{bmatrix}\Delta x\\ \Delta s\end{bmatrix}$
$-\begin{bmatrix}C\Delta x\\ G\Delta x+\Delta s\end{bmatrix}^{T}\begin{bmatrix}\eta(C\Delta x)\\ \eta(G\Delta x+\Delta s)\end{bmatrix}$
$=$ $-\begin{bmatrix}\Delta x^{T}&\Delta s^{T}\end{bmatrix}\begin{bmatrix}P&0\\ 0&S^{-1}Z\end{bmatrix}\begin{bmatrix}\Delta x\\ \Delta s\end{bmatrix}$
$-\eta(\lVert C\Delta x\rVert^{2}+\lVert G\Delta x+\Delta s\rVert^{2})<0$
Thus, the Augmented Barrier-Lagrangian $\mathcal{A}(\cdot,\cdot,y,z;\mu,\eta)$ can be used as the merit function for a line search over the primal variables $(x,s)$.
@@
$\begin{bmatrix}P&C^{T}&G^{T}\\ C&-\frac{1}{\eta}I&0\\ G&0&-(W+\frac{1}{\eta}I)\end{bmatrix}\begin{bmatrix}\Delta x\\ \Delta y\\ \Delta z\end{bmatrix}$ (3)
$=$ $-\begin{bmatrix}\nabla_{x}\mathcal{L}(x,s,y,z;\mu,\eta)\\ c(x)\\ g(x)+\mu Z^{-1}e\end{bmatrix}$,
where $W=Z^{-1}S$.
@@ II-C Applying the regularized IPM to optimal control
$\begin{bmatrix}P&C^{T}&G^{T}\\ C&-\frac{1}{\eta}I&0\\ G&0&-W-\frac{1}{\eta}I\end{bmatrix}\begin{bmatrix}\Delta x\\ \Delta y\\ \Delta z\end{bmatrix}=-\begin{bmatrix}r_{x}\\ r_{y}\\ r_{z}\end{bmatrix}$,
where $\eta>0$ and $W$ is a diagonal matrix with positive elements.
Note that $\Delta z$ can be eliminated via $\Delta z=(W+\frac{1}{\eta}I)^{-1}(G\Delta x+r_{z})$, resulting in
$\begin{bmatrix}P+G^{T}(W+\frac{1}{\eta}I)^{-1}G&C^{T}\\ C&-\frac{1}{\eta}I\end{bmatrix}\begin{bmatrix}\Delta x\\ \Delta y\end{bmatrix}$
$=$ $-\begin{bmatrix}r_{x}+G^{T}(W+\frac{1}{\eta}I)^{-1}r_{z}\\ r_{y}\end{bmatrix}$.
Any component of $\Delta y$ corresponding to an equality constraint other than the dynamics can be eliminated in the same fashion. Importantly, note that the only constraints that have cross-stage dependencies are the dynamics, so these eliminations preserve the stagewise nature of the problem.
@@ II-D Dual-Regularized LQR
Moreover, the matrices $P_{i}$ are expected to be symmetric and positive semi-definite; the matrices $R_{i}$ are expected to be symmetric and positive definite; the scalars $\delta_{0},\ldots,\delta_{N}$ are expected to be positive.
@@
$+\frac{1}{2}x_{N}^{T}Q_{N}x_{N}+q_{N}^{T}x_{N}+y_{0}^{T}(c_{0}-x_{0})-\sum\limits_{i=0}^{N}\frac{\delta_{i}}{2}y_{i}^{T}y_{i}$
@@ IV Sequential Algorithm
If $\delta>0$ and $f(y)=k^{T}y-\frac{\delta}{2}y^{T}y$, then $\max\limits_{y}f(y)=f(\frac{k}{\delta})=\frac{1}{2\delta}\lVert k\rVert^{2}$.
@@
$\nabla f(y)=k-\delta y=0\Rightarrow y=\frac{k}{\delta}$. Moreover, $f(\frac{k}{\delta})=k^{T}\frac{k}{\delta}-\frac{\delta}{2}\frac{k^{T}k}{\delta^{2}}=\frac{k^{T}k}{\delta}-\frac{k^{T}k}{2\delta}=\frac{1}{2\delta}\lVert k\rVert^{2}$. ∎
@@
If $P$ is symmetric and positive semi-definite and $\delta\geq 0$,
$I-(I+\delta P)^{-1}=(I+\delta P)^{-1}\delta P=\delta P(I+\delta P)^{-1}.$
@@
$(I+\delta P)^{-1}\delta P$ $=(I+\delta P)^{-1}(I+\delta P)-(I+\delta P)^{-1}$
$=I-(I+\delta P)^{-1}.$
Similarly,
$\delta P(I+\delta P)^{-1}$ $=(I+\delta P)(I+\delta P)^{-1}-(I+\delta P)^{-1}$
$=I-(I+\delta P)^{-1}.$
@@
If $P$ is positive semi-definite, $\delta>0$, and $f(x)=\frac{1}{2}x^{T}Px+p^{T}x+\frac{1}{2\delta}\lVert c-x\rVert^{2}$, then
$\min\limits_{x}f(x)=$ $f\left((I+\delta P)^{-1}(c-\delta p)\right)$
$=$ $\frac{1}{2}c^{T}\left((I+\delta P)^{-1}P\right)c+p^{T}c$
$-\delta p^{T}(I+\delta P)^{-1}Pc-\frac{\delta}{2}p^{T}(I+\delta P)^{-1}p.$
@@
$\nabla f(x)=Px+p-\frac{c-x}{\delta}=0$
$\Rightarrow$ $(I+\delta P)x=c-\delta p$
$\Rightarrow$ $x=(I+\delta P)^{-1}(c-\delta p).$
Since we can also write
$f(x)=\frac{1}{\delta}\left(\frac{1}{2}x^{T}(I+\delta P)x+(\delta p-c)^{T}x+\frac{1}{2}c^{T}c\right),$
and noting that we use lemma 3 below, it follows that
$f\left((I+\delta P)^{-1}(c-\delta p)\right)$
$=$ $\frac{1}{\delta}\left(\frac{1}{2}(c-\delta p)^{T}(I+\delta P)^{-1}(c-\delta p)\right$.
[…]
$=$ $\frac{1}{2\delta}\left(c^{T}c-(c-\delta p)(I+\delta P)^{-1}(c-\delta p)\right)$
$=$ $\frac{1}{\delta}\left(\frac{1}{2}c^{T}\left(I-(I+\delta P)^{-1}\right)c+\delta p^{T}(I+\delta P)^{-1}c\right$.
$\left.-\frac{\delta^{2}}{2}p^{T}(I+\delta P)^{-1}p\right)$
$=$ $\frac{1}{\delta}\left(\frac{1}{2}c^{T}\left((I+\delta P)^{-1}\delta P\right)c\right$.
$\left.+\delta p^{T}(I-\delta(I+\delta P)^{-1}P)c-\frac{\delta^{2}}{2}p^{T}(I+\delta P)^{-1}p\right)$
$=$ $\frac{1}{2}c^{T}\left((I+\delta P)^{-1}P\right)c+p^{T}c$
$-\delta p^{T}(I+\delta P)^{-1}Pc-\frac{\delta}{2}p^{T}(I+\delta P)^{-1}p.$
@@
$-\sum\limits_{i=k}^{N-1}\frac{\delta_{i+1}}{2}y_{i+1}^{T}y_{i+1}.$
@@
$\min\limits_{x_{k}}V_{k}(x_{k})=\min\limits_{x_{k}}\left(\frac{1}{2}x_{k}^{T}P_{k}x_{k}+p_{k}^{T}x_{k}\right)+V_{k}(0),$
where $P_{k}$ is symmetric and positive semi-definite.
Moreover, there exist $K_{k},k_{k}$ such that eq. 7 is optimal at
$u_{k}=$ $K_{k}x_{k}+k_{k}$
$x_{0}=$ $(I+\delta_{0}P_{0})^{-1}(c_{0}-\delta_{0}p_{0})$
@@
$+x_{k}^{T}M_{k}u_{k}+\frac{1}{2}u_{k}^{T}R_{k}u_{k}-\frac{\delta_{k+1}}{2}y_{k+1}^{T}y_{k+1}$
@@
$\max\limits_{y_{k+1}}y_{k+1}^{T}\left(A_{i}x_{i}+B_{i}u_{i}+c_{k+1}-x_{k+1}\right)-\frac{\delta_{k+1}}{2}y_{k+1}^{T}y_{k+1}$
$=\frac{1}{2\delta_{k+1}}\lVert A_{i}x_{i}+B_{i}u_{i}+c_{k+1}-x_{k+1}\rVert^{2},$
@@
$+\frac{1}{2\delta_{k+1}}\lVert A_{k}x_{k}+B_{k}u_{k}+c_{k+1}-x_{k+1}\rVert^{2}$
@@
$+\frac{1}{2\delta_{k+1}}\lVert A_{k}x_{k}+B_{k}u_{k}+c_{k+1}-x_{k+1}\rVert^{2}+\text{const}.$
@@
$\min\limits_{x_{k+1}}$ $\frac{1}{2}x_{k+1}^{T}P_{k+1}x_{k+1}+p_{k+1}^{T}x_{k+1}$
$+\frac{1}{2\delta_{k+1}}\lVert A_{k}x_{k}+B_{k}u_{k}+c_{k+1}-x_{k+1}\rVert^{2}.$
Letting
$W_{k+1}$ $=(I+\delta_{k+1}P_{k+1})^{-1}P_{k+1},$
applying lemma 4, and discarding additive constant terms, this becomes
@@
$+(r_{k}+M_{k}^{T}x_{k}+B_{k}^{T}W_{k+1}(A_{k}x_{k}+c_{k+1})$
$+B_{k}^{T}(p_{k+1}-\delta_{k+1}W_{k+1}p_{k+1})^{T}u_{k}$
@@
$p_{k+1}^{T}(A_{k}x_{k}+c_{k+1})-\delta_{k+1}p_{k+1}(W_{k+1}A_{k}x_{k}+c_{k+1}),$ (14)
which do not depent on $u_{k}$. We will pick these back up later.
@@
$\max\limits_{y_{0}}\min\limits_{x_{0}}V_{0}(x_{0})+y_{0}^{T}(x_{0}-c_{0})-\frac{\delta_{0}}{2}y_{0}^{T}y_{0},$
@@
$\min\limits_{x_{0}}V_{0}(x_{0})+\frac{1}{2\delta_{0}}\lVert c_{0}-x_{0}\rVert^{2}$
$=\frac{1}{\delta_{0}}\left(\min\limits_{x_{0}}\frac{1}{2}x_{0}^{T}(I+\delta_{0}P_{0})x_{0}+(\delta_{0}p_{0}-c_{0})^{T}x_{0}+c_{0}^{T}c_{0}\right)$
$+V_{0}(0),$
@@
$x_{0}=(I+\delta_{0}P_{0})^{-1}(c_{0}-\delta_{0}p_{0}).$ (20)
@@
Finally, we can re-write eq. 19 and eq. 10 to remove any division by $\delta_{0},\ldots,\delta_{N}$, in the interest of improving numerical stability when $\delta_{0},\ldots,\delta_{N}\rightarrow 0$ as well as extending this method to the case of $\delta_{0},\ldots,\delta_{N}=0$ (recovering the standard backward and forward passses of the standard LQR algorithm).
@@
$(I+\delta_{0}P_{0})x_{0}=-(\delta_{0}p_{0}-c_{0})\Leftrightarrow\frac{1}{\delta_{0}}(c_{0}-x_{0})=P_{0}x_{0}+p_{0},$

### v5

@@ II-B The Regularized Interior Point Method
If $M$ is symmetric and positive-definite, the squared $M$-norm is defined as $\lVert x\rVert_{M}^{2}=x^{T}Mx$.
Definition 4
@@
$\mathcal{A}(x,s,y,z;\mu,\Delta_{C},\Delta_{G})$
$=$ $\mathcal{L}(x,s,y,z;\mu)+\frac{1}{2}(\lVert c(x)\rVert_{\Delta_{C}}^{2}+\lVert g(x)+s\rVert_{\Delta_{G}}^{2}),$
where $\Delta_{C}$ and $\Delta_{G}$ are symmetric positive-definite matrices.
Frequently, $\Delta_{C}$ and $\Delta_{G}$ are set to $\eta I$, where $\eta>0$. However, having separate penalty parameters for different constraints may promote numerical stability and faster convergence.
We call $x,s$ the primal variables and $y,z$ the dual variables. As usual, we require that $s,z>0$ always holds.
@@
$\begin{bmatrix}P&0&C^{T}&G^{T}\\ 0&W^{-1}&0&I\\ C&0&-\Delta_{C}^{-1}&0\\ G&I&0&-\Delta_{G}^{-1}\end{bmatrix}\begin{bmatrix}\Delta x\\ \Delta s\\ \Delta y\\ \Delta z\end{bmatrix}$$\nabla\mathcal{L}$, (1)
where $\nabla\mathcal{L}$ is evaluated at $(x,s,y,z;\mu)$, $P$ is a positive-definite approximation of $\nabla^{2}_{xx}\mathcal{L}(x,s,y,z;\mu)$, $C=J(c)(x)$, $G=J(g)(x)$, $S,Z$ denote the diagonal matrices with entries $s,z$ (respectively), and $W=Z^{-1}S$ (or any symmetric positive-definite approximation thereof).
Below, we will show that primal search direction determined by the regularized interior point method is guaranteed to be a descent direction of $\mathcal{A}(\cdot,\cdot,y,z;\mu,\Delta_{C},\Delta_{G})$ at $(x,s)$ unless the primal variables have already converged.
@@
$\begin{bmatrix}P&0&C^{T}&G^{T}\\ 0&W^{-1}&0&I\\ C&0&-\Delta_{C}^{-1}&0\\ G&I&0&-\Delta_{G}^{-1}\end{bmatrix}\begin{bmatrix}\Delta x\\ \Delta s\\ \Delta y-\Delta_{C}c(x)\\ \Delta z-\Delta_{G}(g(x)+s)\end{bmatrix}$ (2)
$=$ $-\begin{bmatrix}\nabla_{x}\mathcal{A}(x,s,y,z;\mu,\Delta_{C},\Delta_{G})\\ \nabla_{s}\mathcal{A}(x,s,y,z;\mu,\Delta_{C},\Delta_{G})\\ 0\\ 0\end{bmatrix}$
@@
When $\Delta x\neq 0$ or $\Delta s\neq 0$, $D(\mathcal{A}(\cdot,\cdot,y,z;\mu,\Delta_{C},\Delta_{G});(\Delta x,\Delta s))(x,s)<0$.
@@
$D(\mathcal{A}(\cdot,\cdot,y,z;\mu,\Delta_{C},\Delta_{G});(\Delta x,\Delta s))(x,s)$
$=$ $\begin{bmatrix}\Delta x^{T}&\Delta s^{T}\end{bmatrix}\begin{bmatrix}\nabla_{x}\mathcal{A}(x,s,y,z;\mu,\Delta_{C},\Delta_{G})\\ \nabla_{s}\mathcal{A}(x,s,y,z;\mu,\Delta_{C},\Delta_{G})\\ \end{bmatrix}$
$=$ $\begin{bmatrix}\Delta x^{T}&\Delta s^{T}&0&0\end{bmatrix}\begin{bmatrix}\nabla_{x}\mathcal{A}(x,s,y,z;\mu,\Delta_{C},\Delta_{G})\\ \nabla_{s}\mathcal{A}(x,s,y,z;\mu,\Delta_{C},\Delta_{G})\\ 0\\ 0\end{bmatrix}$
$=$ $-\begin{bmatrix}\Delta x^{T}&\Delta s^{T}&0&0\end{bmatrix}\begin{bmatrix}P&0&C^{T}&G^{T}\\ 0&W^{-1}&0&I\\ C&0&-\Delta_{C}^{-1}&0\\ G&I&0&-\Delta_{G}^{-1}\end{bmatrix}$
$\begin{bmatrix}\Delta x\\ \Delta s\\ \Delta y-\Delta_{C}c(x)\\ \Delta z-\Delta_{G}(g(x)+s)\end{bmatrix}$
$=$ $-\begin{bmatrix}\Delta x^{T}&\Delta s^{T}\end{bmatrix}\begin{bmatrix}P&0\\ 0&W^{-1}\end{bmatrix}\begin{bmatrix}\Delta x\\ \Delta s\end{bmatrix}$
$-\begin{bmatrix}\Delta x^{T}&\Delta s^{T}\end{bmatrix}\begin{bmatrix}C^{T}&G^{T}\\ 0&I\end{bmatrix}\begin{bmatrix}\Delta y-\Delta_{C}c(x)\\ \Delta z-\Delta_{G}(g(x)+s)\end{bmatrix}$
$=$ $-\begin{bmatrix}\Delta x^{T}&\Delta s^{T}\end{bmatrix}\begin{bmatrix}P&0\\ 0&W^{-1}\end{bmatrix}\begin{bmatrix}\Delta x\\ \Delta s\end{bmatrix}$
$-\begin{bmatrix}C\Delta x\\ G\Delta x+\Delta s\end{bmatrix}^{T}\begin{bmatrix}\Delta_{C}(C\Delta x)\\ \Delta_{G}(G\Delta x+\Delta s)\end{bmatrix}$
$=$ $-\lVert\Delta x\rVert_{P}^{2}-\lVert\Delta s\rVert_{W^{-1}}^{2}-\lVert C\Delta x\rVert_{\Delta_{C}}^{2}-\lVert G\Delta x+\Delta s\rVert_{\Delta_{G}}^{2}$
$<$ $0$
Thus, $\mathcal{A}(\cdot,\cdot,y,z;\mu,\Delta_{C},\Delta_{G})$ (i.e. the Augmented Barrier-Lagrangian) can be used as the merit function for a line search over the primal variables $(x,s)$.
@@
$\begin{bmatrix}P&C^{T}&G^{T}\\ C&-\Delta_{C}^{-1}&0\\ G&0&-(W+\Delta_{G}^{-1})\end{bmatrix}\begin{bmatrix}\Delta x\\ \Delta y\\ \Delta z\end{bmatrix}$ (3)
$=$ $-\begin{bmatrix}\nabla_{x}\mathcal{L}(x,s,y,z;\mu)\\ c(x)\\ g(x)+s+W(\mu S^{-1}e-z)\end{bmatrix}$.
@@ II-C Applying the regularized IPM to optimal control
$\begin{bmatrix}P&C^{T}&G^{T}\\ C&-\Delta_{C}^{-1}&0\\ G&0&-W-\Delta_{G}^{-1}\end{bmatrix}\begin{bmatrix}\Delta x\\ \Delta y\\ \Delta z\end{bmatrix}=-\begin{bmatrix}r_{x}\\ r_{y}\\ r_{z}\end{bmatrix}$.
Given that $\Delta z=(W+\Delta_{G}^{-1})^{-1}(G\Delta x+r_{z})$, we can also eliminate $\Delta z$, resulting in
$\begin{bmatrix}P+G^{T}(W+\Delta_{G}^{-1})^{-1}G&C^{T}\\ C&-\Delta_{C}^{-1}I\end{bmatrix}\begin{bmatrix}\Delta x\\ \Delta y\end{bmatrix}$
$=$ $-\begin{bmatrix}r_{x}+G^{T}(W+\Delta_{G}^{-1})^{-1}r_{z}\\ r_{y}\end{bmatrix}$.
Any component of $\Delta y$ corresponding to an equality constraint other than the dynamics can be eliminated in the same fashion. Importantly, note that the only constraints that have cross-stage dependencies are the dynamics, so these eliminations preserve the stagewise nature of the problem, provided that $W,\Delta_{C},\Delta_{G}$ are reasonably chosen to not introduce cross-stage dependencies; note that $W,\Delta_{C},\Delta_{G}$ are typically diagonal, in which case this requirement is met.
@@ II-D Dual-Regularized LQR
Moreover, the matrices $P_{i}$ are expected to be symmetric and positive semi-definite, and the matrices $R_{i}$ and $\Delta_{i}$ are expected to be symmetric and positive definite. Typically, the matrices $\Delta_{i}$ would also be diagonal, but we do not impose this requirement.
@@
$+\frac{1}{2}x_{N}^{T}Q_{N}x_{N}+q_{N}^{T}x_{N}-\sum\limits_{i=0}^{N}\frac{1}{2}y_{i}^{T}\Delta_{i}y_{i}+y_{0}^{T}(c_{0}-x_{0})$
@@ IV Sequential Algorithm
If $M$ is symmetric and positive-definite and $f(y)=k^{T}y-\frac{1}{2}y^{T}My$, then $\max\limits_{y}f(y)=f(M^{-1}k)=\frac{1}{2}\lVert k\rVert_{M^{-1}}^{2}$.
@@
$\nabla f(y)=k-My=0\Rightarrow y=M^{-1}k$. Moreover, $f(M^{-1}k)=k^{T}M^{-1}k-\frac{1}{2}k^{T}M^{-1}k=\frac{1}{2}\lVert k\rVert_{M^{-1}}^{2}$. ∎
@@
If $P$ and $M$ are symmetric and positive semi-definite,
$I-(I+MP)^{-1}$ $=MP(I+MP)^{-1}=(I+MP)^{-1}MP,$
$(P+M^{-1})^{-1}$ $=(I+MP)^{-1}M=M(I+PM)^{-1},$
$(I+PM)^{-1}P$ $=P(I+MP)^{-1}.$
@@
$MP(I+MP)^{-1}$ $=(I+MP)(I+MP)^{-1}-(I+MP)^{-1}$
$=I-(I+MP)^{-1}$
$(I+MP)^{-1}MP$ $=(I+MP)^{-1}(I+MP)-(I+MP)^{-1}$
$=I-(I+MP)^{-1}$
$(P+M^{-1})^{-1}$ $=(M^{-1}(I+MP))^{-1}=(I+MP)^{-1}M$
$(P+M^{-1})^{-1}$ $=((I+PM)M^{-1})^{-1}=M(I+PM)^{-1}$
$(I+PM)^{-1}P$ $=((P+M^{-1})M)^{-1}P$
$=M^{-1}(P+M^{-1})^{-1}P$
$=M^{-1}(I+MP)^{-1}MP$
$=M^{-1}MP(I+MP)^{-1}$
$=P(I+MP)^{-1}.$
@@
If $P$ is symmetric and positive semi-definite, $M$ is symmetric and positive definite, and $f(x)=\frac{1}{2}x^{T}Px+p^{T}x+\frac{1}{2}\lVert c-x\rVert_{M^{-1}}^{2}$, then
$\min\limits_{x}f(x)=$ $f\left((I+MP)^{-1}(c-Mp)\right)$
$=$ $\frac{1}{2}c^{T}P(I+MP)^{-1}c-\frac{1}{2}p^{T}(I+MP)^{-1}Mp$
$+$ $p^{T}c-p^{T}MP(I+MP)^{-1}c.$
@@
Noting that
$f(x)=\frac{1}{2}x^{T}(P+M^{-1})x+(p-M^{-1}c)^{T}x+\frac{1}{2}c^{T}M^{-1}c,$
it follows that
$\nabla f(x)=(P+M^{-1})x+p-M^{-1}c=0$
$\Rightarrow$ $(P+M^{-1})x=M^{-1}c-p$
$\Rightarrow$ $x=(P+M^{-1})^{-1}(M^{-1}c-p).$
Moreover, employing lemma 3,
$f\left((P+M^{-1})^{-1}(M^{-1}c-p)\right)$
$=$ $\frac{1}{2}(M^{-1}c-p)^{T}(P+M^{-1})^{-1}(M^{-1}c-p)$
$+$ $(p-M^{-1}c)(P+M^{-1})^{-1}(M^{-1}c-p)+\frac{1}{2}c^{T}M^{-1}c$
$=$ $-\frac{1}{2}(M^{-1}c-p)^{T}(P+M^{-1})^{-1}(M^{-1}c-p)+\frac{1}{2}c^{T}M^{-1}c$
$=$ $\frac{1}{2}c^{T}(M^{-1}-M^{-1}(P+M^{-1})^{-1}M^{-1})c$
- $\frac{1}{2}p^{T}(P+M^{-1})^{-1}p+p^{T}((P+M^{-1})^{-1}M^{-1})c$
$=$ $\frac{1}{2}c^{T}M^{-1}(I-(I+MP)^{-1})c$
- $\frac{1}{2}p^{T}(I+MP)^{-1}Mp+p^{T}(I+MP)^{-1}c$
$=$ $\frac{1}{2}c^{T}P(I+MP)^{-1}c-\frac{1}{2}p^{T}(I+MP)^{-1}Mp$
$+$ $p^{T}c-p^{T}(I-(I+MP)^{-1})c$
$=$ $\frac{1}{2}c^{T}P(I+MP)^{-1}c-\frac{1}{2}p^{T}(I+MP)^{-1}Mp$
$+$ $p^{T}c-p^{T}(MP(I+MP)^{-1})c.$
Finally, note that
$(P+M^{-1})^{-1}(M^{-1}c-p)$ $=(I+MP)^{-1}M(M^{-1}c-p)$
$=(I+MP)^{-1}(c-Mp).$
Definition 7
@@
$-\sum\limits_{i=k}^{N-1}\frac{1}{2}y_{i+1}^{T}\Delta_{i+1}y_{i+1}.$
@@
$V_{k}(x_{k})=\frac{1}{2}x_{k}^{T}P_{k}x_{k}+p_{k}^{T}x_{k}+V_{k}(0),$
where $P_{k}$ is symmetric and positive semi-definite, as well as $K_{k},k_{k}$ such that eq. 7 is optimal at
$u_{k}=$ $K_{k}x_{k}+k_{k},$
$x_{0}=$ $(I+\Delta_{0}P_{0})^{-1}(c_{0}-\Delta_{0}p_{0}),$
@@
$+x_{k}^{T}M_{k}u_{k}+\frac{1}{2}u_{k}^{T}R_{k}u_{k}-\frac{1}{2}y_{k+1}^{T}\Delta_{k+1}y_{k+1}$
@@
$\max\limits_{y_{k+1}}y_{k+1}^{T}\left(A_{i}x_{i}+B_{i}u_{i}+c_{k+1}-x_{k+1}\right)-\frac{1}{2}y_{k+1}^{T}\Delta_{k+1}y_{k+1}$
$=\frac{1}{2}\lVert A_{i}x_{i}+B_{i}u_{i}+c_{k+1}-x_{k+1}\rVert_{\Delta_{k+1}^{-1}}^{2},$
@@
$+\frac{1}{2}\lVert A_{k}x_{k}+B_{k}u_{k}+c_{k+1}-x_{k+1}\rVert_{\Delta_{k+1}^{-1}}^{2}$
@@
$+\frac{1}{2}\lVert A_{k}x_{k}+B_{k}u_{k}+c_{k+1}-x_{k+1}\rVert_{\Delta_{k+1}^{-1}}^{2}+\text{const}.$
@@
$\min\limits_{x_{k+1}}$ $\frac{1}{2}x_{k+1}^{T}P_{k+1}x_{k+1}+p_{k+1}^{T}x_{k+1}$ (12)
$+\frac{1}{2}\lVert A_{k}x_{k}+B_{k}u_{k}+c_{k+1}-x_{k+1}\rVert_{\Delta_{k+1}^{-1}}^{2}.$
Letting $W_{k+1}=P_{k+1}(I+\Delta_{k+1}P_{k+1})^{-1}$, it follows from lemma 3 that $W_{k+1}=W_{k+1}^{T}$.
Applying lemma 4, and discarding additive constant terms, eq. 12 becomes
@@
[…]
$\left.\left.+p_{k+1}-W_{k+1}\Delta_{k+1}p_{k+1}\right)\right)^{T}u_{k}$
@@
$p_{k+1}^{T}(A_{k}x_{k}+c_{k+1})-p_{k+1}^{T}\Delta_{k+1}W_{k+1}(A_{k}x_{k}+c_{k+1}),$ (15)
which do not depend on $u_{k}$. We will pick these back up later.
@@
$\max\limits_{y_{0}}\min\limits_{x_{0}}V_{0}(x_{0})+y_{0}^{T}(x_{0}-c_{0})-\frac{1}{2}y_{0}^{T}\Delta_{0}y_{0},$
@@
$\min\limits_{x_{0}}$ $V_{0}(x_{0})+\frac{1}{2}\lVert c_{0}-x_{0}\rVert_{\Delta_{0}^{-1}}^{2}$
$=\min\limits_{x_{0}}$ $\frac{1}{2}x_{0}^{T}(P_{0}+\Delta_{0}^{-1})x_{0}+(p_{0}-\Delta_{0}^{-1}c_{0})^{T}x_{0}$
$+\frac{1}{2}\lVert c_{0}\rVert_{\Delta_{0}}^{2}+V_{0}(0),$
@@
$x_{0}$ $=(P_{0}+\Delta_{0}^{-1})^{-1}(\Delta_{0}^{-1}c_{0}-p_{0})$ (21)
$=(I+\Delta_{0}P_{0})^{-1}\Delta_{0}(\Delta_{0}^{-1}c_{0}-p_{0})$
$=(I+\Delta_{0}P_{0})^{-1}(c_{0}-\Delta_{0}p_{0}).$
@@
Finally, we can re-write eq. 20 and eq. 10 to remove any usage of $\Delta_{0}^{-1},\ldots,\Delta_{N}^{-1}$, in the interest of improving numerical stability when $\Delta_{0},\ldots,\Delta_{N}\rightarrow 0$ as well as extending this method to the case of $\Delta_{0},\ldots,\Delta_{N}=0$ (recovering the standard backward and forward passses of the standard LQR algorithm).
@@
$(P_{0}+\Delta_{0}^{-1})x_{0}$ $=\Delta_{0}^{-1}c_{0}-p_{0}$
$\Leftrightarrow\Delta_{0}^{-1}(c_{0}-x_{0})$ $=P_{0}x_{0}+p_{0},$